Suppose that the quantity of a good demanded rises from 90 units to 110 units when the price falls from $1.20 to 80 cents per unit. The own price elasticity of demand for this product is

A)- 0.5.

B)   1.0.
C)   1.5.
D)   0.5
E)   4.0.

If the price elasticity of demand for some good is 2.7, a 2 percent increase in the price results in

A)   a 27% decrease in the quantity demanded.
B)   a 5.4% increase in the quantity demanded.
C)   a 27% decrease in the quantity demanded.
D)   a 5.4% decrease in the quantity demanded.
E)   There is not enough information to answer this question.
Answer #1

1) -0.5

Elasticity = (110-90)(1.2+.80) / (.80-1.2)(90+110) = -40/80 = -0.5

2) D is correct

Elasticity = %change in quantity / %change in price

-2.7 = % change in quantity / 2

% change in quantity = 5.4